# Investigating the heterogenous effects of a massive content moderation intervention via Difference-in-Differences

### v3

@@ Abstract
In today’s online environments, users encounter harm and abuse on a daily basis. Therefore, content moderation is crucial to ensure their safety and well-being. However, the effectiveness of many moderation interventions is still uncertain. Here, we apply a causal inference approach to shed light on the effectiveness of The Great Ban, a massive social media deplatforming intervention. We analyze 53M comments shared by nearly 34K users, providing in-depth results on both the intended and unintended consequences of the ban. Our causal analyses reveal that 15.6% of the moderated users abandoned the platform while the remaining ones decreased their overall toxicity by 4.1%. Nonetheless, a subset of those users increased their toxicity by 70% after the intervention. However, the increases in toxicity did not lead to marked increases in activity or engagement, meaning that the most toxic users had an overall limited impact. Our findings bring to light new insights on the effectiveness of deplatforming moderation interventions. Furthermore, they also contribute to informing future content moderation strategies.
@@
[1]organization=University of Pisa - Department of Information Engineering,addressline=Largo Lazzarino 1, postcode=56122, city=Pisa, country=Italy
@@
[2]organization=University of Pisa - Department of Computer Science,addressline=Largo Bruno Pontecorvo 3, postcode=56122, city=Pisa, country=Italy
[…]
[3]organization=IIT-CNR,addressline=Via Moruzzi 1, postcode=56122, city=Pisa, country=Italy
@@ 1 Introduction
Content moderation is a fundamental component of many online platforms as it prevents the spread of problematic content and behavior, such as hate speech (Giorgi et al., 2024; Kocoń et al., 2021), fake news  (Bazmi et al., 2023; Augenstein et al., 2024), and polarization  (Tao et al., 2025; Trujillo and Cresci, 2023). Content moderation also contributes to the fairness and safety of online users by promoting high ethical standards and fostering the health of online communities (Gillespie, 2018). In practice, administrators enforce platform policies through the application of a wide array of moderation interventions, which can vary in severity (Trujillo et al., 2025). For example, moderators can send short warning messages and use informative labels (Katsaros et al., 2022), or even take the drastic decision to remove content and/or users (Trujillo and Cresci, 2022; Jhaver et al., 2021; Russo et al., 2023). However, despite the increasing reliance on content moderation, there is still little understanding of the effects of many interventions, which hinders our capacity to design and apply effective moderation. For example, recent studies demonstrated that while some interventions had the desired effects, others resulted in heterogeneous (Horta Ribeiro et al., 2021; Trujillo and Cresci, 2023) or even ineffectual outcomes (Dias et al., 2020), or even led to undesirable consequences (Bail et al., 2018; Pennycook et al., 2020). Therefore, assessing the effects of moderation interventions represents an important preliminary step towards developing new and more effective forms of moderation.
@@
The most popular and widely used type of intervention is deplatforming, which involves the removal of content, users, or even entire communities (Horta Ribeiro et al., 2024). Notorious examples are the ban that Donald Trump received in 2021 from Facebook and X (formerly Twitter) (Seeliger and Baum, 2023) and the deplatforming of three particularly toxic influencers from X  (Jhaver et al., 2021). Additionally, X removed accounts involved in coordinated inauthentic behavior (Cima et al., 2024a) and Reddit permanently shut down different communities because of racism, sexism, and hatefulness (Chandrasekharan et al., 2017; Habib et al., 2022). In June 2020, Reddit itself hosted one of the biggest deplatforming campaigns in the history of social media –The Great Ban– which resulted in around 2,000 subreddits being banned due to ongoing spread of toxicity and hate speech. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update_to_our_content_policy/ (accessed: 10/15/2024) Among these are popular communities such as r/The_Donald and r/ChapoTrapHouse. Despite its impact on several communities and users within and outside Reddit, its effects are still little-explored. For example, current research on The Great Ban has mainly investigated the changes in the writing style of the users (Trujillo et al., 2021) without assessing the ban’s effectiveness. Few studies analyzed the changes in toxicity, and those who did focused only on a small set of subreddits (Trujillo and Cresci, 2022). Moreover, the majority of the existing works on The Great Ban are focused on community-level effects, neglecting the individual user-level responses that are instrumental for understanding how effective the ban was in mitigating problematic behaviors (Robertson, 2022; Trujillo and Cresci, 2023). In a previous work, we provided preliminary results on the effectiveness and unintended consequences of The Great Ban (Cima et al., 2024b). Here, we extend our previous analysis by adopting a robust causal approach based on Difference-in-Differences (DiD) to estimate the effects of the intervention. With respect to our previous work we also address an additional research question investigating the impact that toxic users had after the intervention, as explained in the following.
@@ 2 Related Work
In this section, we review and analyze recent literature on the evaluation of moderation interventions, beginning with studies most closely related to our own.
@@ 2.1 Deplatforming
The Great Ban was a significant event that had a major impact on Reddit. However, only a few studies actually looked into its effects. One of them is the study conducted by Trujillo et al. (2021), that investigated the changes in activity and use of language among the 15 most popular subreddits involved in the ban. They discovered that the most active users drastically decreased their level of activity and observed heterogeneous responses to the ban both intra- and inter- subreddits. Here, we build on this study by evaluating the impact of The Great Ban in terms of toxicity, rather than focusing on activity and language. Another body of works focused on assessing the effects of deplatforming in a subset of subreddits affected by the ban or in entirely different subreddits. Chandrasekharan et al. (2017) and Saleem and Ruths (2018) evaluated how The Great Ban impacted two specific subreddits: r/fatpeoplehate and r/coontown. They revealed how a large portion of users abandoned the platform and the remaining ones notably decreased their toxicity levels. However, they also observed that several users migrated to other subreddits and doubled their posting activity (Chandrasekharan et al., 2022). Other studies focused on deplatforming both within and outside Reddit. Horta Ribeiro et al. (2021) studied the migration of Reddit users from banned subreddits to other platforms after a deplatforming intervention. They show how migrated users became drastically less active on the new platforms. This was not the case for a group of users, who instead became much more toxic and radicalized.  Mekacher et al. (2023) analyzed the consequences of deplatforming from Twitter to Gettr. They found that politically polarized users exhibit lower toxicity on fringe platforms, likely due to reduced exposure to interactions with out-group members. Moreover, Cima et al. (2024a) examined large-scale bans implemented by Twitter to combat coordinated inauthentic behaviors. To conclude,  Jhaver et al. (2021) studied the impact of the banishment of three notorious influencers from Twitter. They observed a general decrease in attention towards these influencers and toxicity among their fans. However, some users became much more active and toxic.
@@
In general, these studies highlight how moderation interventions do not always yield the desired results and that different communities and different users can react differently to the same intervention. Our work aims to extend the existing literature by shedding light on the effects of The Great Ban, one of the biggest – yet little-explored – moderation interventions. Our study provides a comprehensive evaluation of the ban’s impact on the 15 most popular subreddits in terms of toxicity, activity, and engagement.
@@
IN-BEFORE OUT-BEFORE OUT-AFTER subreddit subscribers core users comments users comments users comments r/chapotraphouse 159,185 9,295 1,368,874 9,205 3,947,894 8,319 3,157,462 r/the_donald 792,050 4,262 619,434 4,132 2,578,026 3,145 1,434,008 r/darkhumorandmemes 421,506 1,632 35,561 1,617 1,246,399 1,392 689,079 r/consumeproduct 64,937 1,730 60,073 1,719 1,209,933 1,275 594,349 r/gendercritical 64,772 1,091 94,735 1,039 511,173 706 287,877 r/thenewright 41,230 729 5,792 726 600,057 575 308,584 r/soyboys 17,578 596 5,102 594 454,659 432 190,570 r/shitneoconssay 8,701 559 9,178 555 338,218 384 140,619 r/debatealtright 7,381 488 27,814 476 274,600 328 117,281 r/darkjokecentral 185,399 316 3,214 308 307,876 270 179,067 r/wojak 26,816 244 1,666 240 210,249 170 81,142 r/hatecrimehoaxes 20,111 189 775 188 185,379 143 96,457 r/ccj2 11,834 150 9,785 145 101,165 119 63,393 r/imgoingtohellforthis2 47,363 93 376 92 74,664 72 43,018 r/oandaexclusiveforum 2,389 60 1,313 59 48,774 55 35,853 focus group total (unique) 16,828 2,243,692 16,540 8,235,086 13,963 5,592,321 r/askreddit 44,279,043 15,863 3,220,750 15,850 17,014,057 15,576 12,372,977 r/worldnews 34,091,866 10,051 577,777 10,050 12,728,027 9,907 9,517,097 r/politics 8,438,064 9,002 1,862,399 8,998 11,059,478 8,862 8,328,241 baseline group total (unique) 16,955 5,660,926 16,937 18,480,693 16,633 13,378,316
@@
Table 1: Composition of the dataset. Subreddits are sorted by number of core users before the ban. Data in IN-BEFORE refers to user activity within the specified subreddits before the ban occurred. OUT-BEFORE and OUT-AFTER represent user activity outside of the specified subreddits (i.e., in the rest of Reddit), respectively before and after the ban.
@@ 2.2 Soft moderation
Despite deplatforming being one of the most widely adopted moderation interventions, it can be perceived by many users as a threat to free speech since it involves the removal of content and users (Zannettou, 2021). This led to the rise of an alternative category of interventions known as soft interventions. Recently, these have been the focus of several studies. For example, Trujillo and Cresci (2022, 2023) investigated the effects of quarantines and restrictions, which are frequently implemented before community bans on Reddit. In particular, they analyzed the effects of moderation on r/the_donald, revealing a decrease in overall activity and toxicity. However, this came at the expense of increased political polarization and decreased factual accuracy in shared news (Trujillo and Cresci, 2022, 2023).  Chandrasekharan et al. (2022) and Shen and Rosé (2022) also studied the quarantine of r/the_donald. They determined that this intervention was essentially ineffectual and did not produce notable changes in terms of misogynistic and racist comments or users’ engagement and behavior.
@@
Another common example is flagging disputed posts with warning labels, whose effectiveness was evaluated in terms of perceived credibility and user engagement.  Pennycook et al. (2020) revealed that the use of warning labels can enhance the perceived credibility of unlabeled posts, including those containing false information that has not yet been addressed. In addition,  Zannettou (2021) demonstrated how users tend to reply often to tweets containing warning labels in order to further debunk the claims. However, this leads to labeled tweets circulating more than unlabelled ones. To conclude, Katsaros et al. (2022) assessed the effectiveness of showing warning messages to users before they post toxic tweets through an A/B test on Twitter. The test confirmed that this strategy is effective at reducing the number of toxic content on average, although a small fraction of users actually increased their toxicity.
The findings on soft moderation interventions align with the findings on deplatforming, showing that each intervention can produce both desired and undesired effects. Overall, these studies emphasize the necessity for additional research to assess the impact of less-examined interventions.
@@ 3 Dataset
Our dataset [^2] https://doi.org/10.5281/zenodo.14034510 includes 16M comments shared by 16,828 distinct Reddit users who have been active in at least one of the 15 most popular public subreddits that were banned during The Great Ban (Trujillo et al., 2021), as detailed in Table 1. In this context, the term popular refers to the number of daily active users. Even though Reddit administrators initially published an obfuscated list of the most popular banned subreddits, [^3] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 10/15/2024) Trujillo et al. (2021) were able to decipher this list for subreddits with over 2,000 daily active users. A total of 15 subreddits were identified. The composition of our dataset is shown in Figure 1, and the procedure adopted to build it is described below.
@@
Focus group. We began by collecting all comments posted between December 2019 and June 2020 in each of the 15 popular banned subreddits, yielding 8M comments from 194K distinct users. To this end, we used the Pushshift data archive (Baumgartner et al., 2020). The dataset covers 30 weeks (7 months) leading up to The Great Ban, providing a suitable baseline for user activity before the moderation intervention (Trujillo and Cresci, 2023). Notably, we were unable to collect data between May and June 2020 due to some subreddits halting their activity or being banned prior to Reddit’s public announcement of The Great Ban on June 29, 2020, as shown in Figure 1.
@@
Following recent literature (Trujillo and Cresci, 2022; Bouleimen et al., 2023), we identified a representative set of users for the selected subreddits by focusing on core users, that is users who consistently participated in at least one subreddit. Core users were characterized as those who posted at least one comment each month from December 2019 to March 2020.444We excluded April 2020 from this analysis due to the limited number of collected comments, likely because subreddit activity had already slowed or halted in the early days of that month. Moreover, we excluded bots (i.e., clearly automated accounts) by removing all accounts that posted two or more comments at the exact same timestamp (Hurtado et al., 2019). To this end, we implemented a minimum time interval of 1 second between comments, ensuring that we did not unintentionally exclude genuine users. We verified the validity of this approach by manually validating a random sample of 1,000 comments. After these filtering steps, we ended up with 2.2M comments made by 16,828 core users. Henceforth, we denote this subset of our dataset as IN-BEFORE, representing activity within the banned subreddits prior to the ban.
@@
To ensure a fair evaluation of the effects of The Great Ban, we need to match comparable datasets before and after the intervention. Since no activity exists inside the banned subreddits following the intervention, we collected all comments made by the core users outside the 15 banned subreddits during the 7 months before and after the ban as shown in Figure 1. We gathered approximately 13.8M comments from 16,540 distinct users. Data related to user activities before the ban was labeled as OUT-BEFORE, and data after the ban as OUT-AFTER, as shown in Table 1. Estimates of the ban’s effects were derived by comparing the OUT-BEFORE and OUT-AFTER datasets.

### v4

# Investigating the heterogeneous effects of a massive content moderation intervention via Difference-in-Differences

@@ Abstract
In today’s online environments, users encounter harm and abuse on a daily basis. Therefore, content moderation is crucial to ensure their safety and well-being. However, the effectiveness of many moderation interventions is still uncertain. Here, we apply a causal inference approach to shed light on the effectiveness of The Great Ban, a massive social media deplatforming intervention on Reddit. We analyze 53M comments shared by nearly 34K users, providing in-depth results on both the intended and unintended consequences of the ban. Our causal analyses reveal that 15.6% of the moderated users abandoned the platform while the remaining ones decreased their overall toxicity by 4.1%. Nonetheless, a small subset of users exhibited marked increases in both the intensity and volume of toxic behavior, particularly among those whose activity levels changed after the intervention. However, these reactions were not accompanied by greater activity or engagement, suggesting that even the most toxic users maintained a limited overall impact. Our findings bring to light new insights on the effectiveness of deplatforming moderation interventions. Furthermore, they also contribute to informing future content moderation strategies and regulations.
Article published in Online Social Networks and Media (OSNEM). DOI: 10.1016/j.osnem.2025.100320. Please, cite the published version.
@@ 1 Introduction
Content moderation is a fundamental component of many online platforms as it prevents the spread of problematic content and behavior, such as hate speech [25, 40, 62], fake news  [7, 4, 16], and polarization  [53, 57, 21]. Content moderation also contributes to the fairness and safety of online users by promoting high ethical standards and fostering the health of online communities [24]. In practice, administrators enforce platform policies through the application of a wide array of moderation interventions, which can vary in severity [58]. For example, moderators can send short warning messages and use informative labels [37] or even take the drastic decision to remove content and/or users [56, 34, 46]. However, despite the increasing reliance on content moderation, there is still little understanding of the effects of many interventions, which hinders our capacity to design and apply effective moderation. For example, recent studies demonstrated that while some interventions had the desired effects, others resulted in heterogeneous [31, 57] or even ineffectual outcomes [19], or even led to undesirable consequences [5, 43]. Therefore, assessing the effects of moderation interventions represents an important preliminary step towards developing new and more effective forms of moderation.
@@
The most popular and widely used type of intervention is deplatforming, which involves the removal of content, users, or even entire communities [30]. Notorious examples are the ban that Donald Trump received in 2021 from Facebook and X (formerly Twitter) [50] and the deplatforming of three particularly toxic influencers from X  [34]. Additionally, X removed accounts involved in coordinated inauthentic behavior [11] and Reddit permanently shut down different communities because of racism, sexism, and hatefulness [10, 27]. In June 2020, Reddit itself hosted one of the biggest deplatforming campaigns in the history of social media –The Great Ban– which resulted in around 2,000 subreddits being banned due to ongoing spread of toxicity and hate speech. [^1] https://www.reddit.com/r/announcements/comments/hi3oht/update_to_our_content_policy/ (accessed: 10/15/2024) Among these are popular communities such as r/The_Donald and r/ChapoTrapHouse. Despite its impact on several communities and users within and outside Reddit, its effects are still little-explored. For example, current research on The Great Ban has mainly investigated the changes in the writing style of the users [59] without assessing the ban’s effectiveness. Few studies analyzed the changes in toxicity, and those who did focused only on a small set of subreddits [56]. Moreover, the majority of the existing works on The Great Ban are focused on community-level effects, neglecting the individual user-level responses that are instrumental for understanding how effective the ban was in mitigating problematic behaviors [45, 57]. In a previous work, we provided preliminary results on the effectiveness and unintended consequences of The Great Ban [13]. Here, we extend our previous analysis by adopting a robust causal inference approach to estimate the effects of the intervention. Specifically, we employ a Difference-in-Differences (DiD) methodology, which allows us to quantify changes in user behavior while accounting for general platform-wide trends. To do so, we construct a large-scale dataset comprising 53 million comments from nearly 34,000 users, covering the seven months before and after the ban. We define a focus group of users who were active in the banned subreddits and compare their post-ban behavior to that of a baseline group of thematically related but unaffected subreddits. We then analyze changes in toxicity, activity, and engagement to assess both the intended and unintended consequences of the Great Ban. With respect to our previous work we also address an additional research question investigating the impact that toxic users had after the intervention, as explained in the following.
@@ 2 Related Work
We review and analyze recent literature on the evaluation of moderation interventions, beginning with studies most closely related to our own.
@@ 2.1 Deplatforming
The Great Ban was a significant event that had a major impact on Reddit. However, only a few studies actually looked into its effects. One of them is the study conducted by M. Trujillo et al. [59], that investigated the changes in activity and use of language among the 15 most popular subreddits involved in the ban. They discovered that the most active users drastically decreased their level of activity and observed heterogeneous responses to the ban both intra- and inter-subreddits. Here, we build on this study by evaluating the impact of The Great Ban in terms of toxicity, rather than focusing on activity and language. Another body of works focused on assessing the effects of deplatforming in a subset of subreddits affected by the ban or in entirely different subreddits. Chandrasekharan et al. [10] and Saleem and Ruths [47] evaluated how The Great Ban impacted two specific subreddits: r/fatpeoplehate and r/coontown. They revealed how a large portion of users abandoned the platform and the remaining ones notably decreased their toxicity levels. However, they also observed that several users migrated to other subreddits and doubled their posting activity [9]. Other studies focused on deplatforming both within and outside Reddit. Horta Ribeiro et al. [31] studied the migration of Reddit users from banned subreddits to other platforms after a deplatforming intervention. They show how migrated users became drastically less active on the new platforms. This was not the case for a group of users, who instead became much more toxic and radicalized.  Mekacher et al. [42] analyzed the consequences of deplatforming from Twitter to Gettr. They found that politically polarized users exhibit lower toxicity on fringe platforms, likely due to reduced exposure to interactions with out-group members. Moreover, Jhaver et al. [34] studied the impact of the banishment of three notorious influencers from Twitter. They observed a general decrease in attention towards these influencers and toxicity among their fans. However, some users became much more active and toxic. Finally, Cima et al. [11], Schoch et al. [48], and Seckin et al. [49] examined large-scale bans implemented by Twitter to combat coordinated inauthentic behaviors. While these studies did not assess the effects of such bans, they tackled the challenging task of constructing a control group to compare the behavior of banned users (i.e., the treatment group). Their findings underscore the inherent difficulties in identifying suitable control groups for causal analyses in online moderation studies.
@@
In general, the previous studies highlight how moderation interventions do not always yield the desired results and that different communities and different users can react differently to the same intervention. Our work aims to extend the existing literature by shedding light on the effects of The Great Ban, one of the biggest – yet little-explored – moderation interventions. Our study provides a comprehensive evaluation of the ban’s impact on the 15 most popular subreddits in terms of toxicity, activity, and engagement.
@@
IN-BEFORE OUT-BEFORE OUT-AFTER subreddit subscribers core users comments users comments users (%) comments (%) r/chapotraphouse 159,185 9,295 1,368,874 9,205 3,947,894 8,319 (90.37) 3,157,462 (79.98) r/the_donald 792,050 4,262 619,434 4,132 2,578,026 3,145 (76.11) 1,434,008 (55.62) r/darkhumorandmemes 421,506 1,632 35,561 1,617 1,246,399 1,392 (86.09) 689,079 (55.29) r/consumeproduct 64,937 1,730 60,073 1,719 1,209,933 1,275 (74.17) 594,349 (49.12) r/gendercritical 64,772 1,091 94,735 1,039 511,173 706 (67.95) 287,877 (56.32) r/thenewright 41,230 729 5,792 726 600,057 575 (79.20) 308,584 (51.43) r/soyboys 17,578 596 5,102 594 454,659 432 (72.73) 190,570 (41.91) r/shitneoconssay 8,701 559 9,178 555 338,218 384 (69.19) 140,619 (41.58) r/debatealtright 7,381 488 27,814 476 274,600 328 (68.91) 117,281 (42.71) r/darkjokecentral 185,399 316 3,214 308 307,876 270 (87.66) 179,067 (58.16) r/wojak 26,816 244 1,666 240 210,249 170 (70.83) 81,142 (38.59) r/hatecrimehoaxes 20,111 189 775 188 185,379 143 (76.06) 96,457 (52.03) r/ccj2 11,834 150 9,785 145 101,165 119 (82.07) 63,393 (62.66) r/imgoingtohellforthis2 47,363 93 376 92 74,664 72 (78.26) 43,018 (57.62) r/oandaexclusiveforum 2,389 60 1,313 59 48,774 55 (93.22) 35,853 (73.51) focus group total (unique) 16,828 2,243,692 16,540 8,235,086 13,963 (84.42) 5,592,321 (67.91) r/askreddit 44,279,043 15,863 3,220,750 15,850 17,014,057 15,576 (98.27) 12,372,977 (72.72) r/worldnews 34,091,866 10,051 577,777 10,050 12,728,027 9,907 (98.58) 9,517,097 (74.77) r/politics 8,438,064 9,002 1,862,399 8,998 11,059,478 8,862 (98.49) 8,328,241 (75.30) baseline group total (unique) 16,955 5,660,926 16,937 18,480,693 16,633 (98.21) 13,378,316 (72.39)
@@
Table 1: Composition of the dataset. Subreddits are sorted by number of core users before the ban. Data in IN-BEFORE refers to user activity within the specified subreddits before the ban occurred. OUT-BEFORE and OUT-AFTER represent user activity outside of the specified subreddits (i.e., in the rest of Reddit), respectively before and after the ban. The percentages in the OUT-AFTER columns indicate the fraction of users and comments after the ban with respect to those in OUT-BEFORE.
@@ 2.2 Soft moderation
Despite deplatforming being one of the most widely adopted moderation interventions, it can be perceived by many users as a threat to free speech, as it involves the removal of content and users [63]. This concern has led to the rise of an alternative category of interventions known as soft interventions, which attempt to mitigate harmful behavior without outright banning users or communities. Recently, these strategies have been the focus of several studies. For example, Trujillo and Cresci [56, 57] investigated the effects of quarantines and restrictions, which are moderation measures frequently implemented before community bans on Reddit. Quarantines limit a subreddit’s visibility by removing it from search results and requiring users to opt in before viewing its content, while restrictions can include measures such as disabling specific posting privileges or preventing the subreddit from generating revenue. These interventions serve as intermediate steps to curb harmful behavior before resorting to a full ban. In particular, they analyzed the effects of moderation on r/the_donald, revealing a decrease in overall activity and toxicity. However, this came at the expense of increased political polarization and decreased factual accuracy in shared news [56, 57]. Chandrasekharan et al. [9] and Shen and Rosé [52] also studied the quarantine of r/the_donald, finding that this intervention was largely ineffectual and did not produce notable changes in misogynistic and racist comments or user engagement.
@@
Another common example is flagging disputed posts with warning labels, whose effectiveness has been studied in terms of perceived credibility and user engagement. Pennycook et al. [43] found that warning labels can increase the perceived credibility of unlabeled posts, potentially amplifying misinformation that has not been flagged. In addition, Zannettou [63] demonstrated that tweets with warning labels often receive higher engagement, as users attempt to debunk the claims. This paradoxical effect suggests that soft moderation can sometimes lead to unintended amplification, rather than suppression, of harmful content. Similarly, Katsaros et al. [37] assessed the effectiveness of preemptive warning messages shown to users before they post toxic tweets. Their A/B test on Twitter confirmed that this strategy was generally effective at reducing toxic content, although a small fraction of users actually increased their toxicity in response to the warnings.
While soft moderation offers a less disruptive alternative to deplatforming, these studies highlight its mixed effectiveness. Unlike hard moderation, soft interventions allow communities to adapt to restrictions rather than being eliminated outright. However, they also risk reinforcing echo chambers by isolating communities without fully dismantling them [31]. For instance, Chandrasekharan et al. [9] found that quarantined subreddits often develop stronger internal cohesion and more extreme discourse, which can counteract the intended moderation effects. The trade-offs between soft and hard moderation suggest that no single approach is universally effective [17]. Deplatforming is more decisive, removing toxic actors from a platform but potentially leading to migration effects [29]. Soft moderation, on the other hand, allows for gradual adaptation, but it can be circumvented, may have limited effectiveness, and can sometimes inadvertently strengthen harmful communities [9]. These nuances highlight the need for hybrid moderation strategies, where soft interventions serve as preliminary measures before escalating to deplatforming when necessary [39].
Overall, these studies emphasize that both soft and hard moderation interventions come with inherent trade-offs, and additional research is needed to assess how different strategies interact and whether a combination of approaches can yield better moderation outcomes.
@@ 3 Dataset
Our dataset [^2] https://doi.org/10.5281/zenodo.14034510 includes 16M comments shared by 16,828 distinct Reddit users who have been active in at least one of the 15 most popular public subreddits that were banned during The Great Ban [59], as detailed in Table 1. In this context, the term popular refers to the number of daily active users. Even though Reddit administrators initially published an obfuscated list of the most popular banned subreddits, [^3] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 10/15/2024) M. Trujillo et al. [59] were able to decipher this list for subreddits with over 2,000 daily active users. A total of 15 subreddits were identified. The composition of our dataset is shown in Figure 1, and the procedure adopted to build it is described below.
3.1 Focus group
@@
We began by collecting all comments posted between December 2019 and June 2020 in each of the 15 popular banned subreddits, yielding 8M comments from 194K distinct users. To this end, we used the Pushshift data archive [6]. The dataset covers 30 weeks (7 months) leading up to The Great Ban, providing a suitable baseline for user activity before the moderation intervention [57]. Notably, we were unable to collect data between May and June 2020 due to some subreddits halting their activity or being banned prior to Reddit’s public announcement of The Great Ban on June 29, 2020, as shown in Figure 1.
@@
Following recent literature [56, 8], we identified a representative set of users for the selected subreddits by focusing on core users, that is users who consistently participated in at least one subreddit. Core users were characterized as those who posted at least one comment each month from December 2019 to March 2020.444We excluded April 2020 from this analysis due to the limited number of collected comments, likely because subreddit activity had already slowed or halted in the early days of that month. Moreover, we excluded bots (i.e., clearly automated accounts) by removing all accounts that posted two or more comments at the exact same timestamp [32]. To this end, we implemented a minimum time interval of 1 second between comments, ensuring that we did not unintentionally exclude genuine users. This process led to the identification of 741 automated accounts among the 17,569 core users (4.22%), which we discarded. We verified the validity of this approach by manually validating a random sample of 1,000 comments. The automated accounts exhibited repetitive posting patterns and often engaged in behaviors characteristic of information dissemination bots, such as mass-posting external links or reposting the same messages across multiple threads. After these filtering steps, we ended up with 2.2M comments made by 16,828 core users. Henceforth, we denote this subset of our dataset as IN-BEFORE, representing activity within the banned subreddits prior to the ban.
@@
To ensure a fair evaluation of the effects of The Great Ban, we need to match comparable datasets before and after the intervention. Since no activity exists inside the banned subreddits following the intervention, we collected all comments made by the core users outside the 15 banned subreddits during the 7 months before and after the ban as shown in Figure 1. All user-level analyses are based on immutable user IDs, ensuring that user activity is accurately tracked across the pre- and post-ban periods. We gathered approximately 13.8M comments from 16,540 distinct users. Data related to user activities before the ban was labeled as OUT-BEFORE, and data after the ban as OUT-AFTER, as shown in Table 1. Estimates of the ban’s effects were derived by comparing the OUT-BEFORE and OUT-AFTER datasets.